# Can Large Reasoning Models Self-Train?

### v1

@@ B.1 RLOO Algorithm
Sampling
For all experiments, we kept the generation temperature to 1.0, top_k and top_p to -1 for both RL rollouts and evaluation. We cut off maximum prompt length at 1024 and maximum response length to 3072, since Qwen2.5-Math-7B models support a maximum context window of 4096.
@@ B.2 GPU Infrastructure
All experiments in this work were conducted using either a single node consisting of 8 NVIDIA H200 GPUs (141 GB of GPU memory per GPU) or a single node consisting of 4 NVIDIA GH200 GPUs (96 GB of GPU memory per GPU). All experiments can be replicated in single-node training, and we did not, in fact, utilize multinode training. In total, this work consumed $\sim$8000 GPU hours (including preliminary studies and failed runs). All the final results listed in this paper can be replicated within 1000 H200 GPU hours.
B.3 Code and Dataset Release
@@
    The test datasets compiled together: https://huggingface.co/datasets/ftajwar/srt_test_dataset
@@
The easy DAPO subsets are the most important datasets, since they allow quick reproduction of our results in Section 5.3. Finally, our code is available at: https://github.com/tajwarfahim/srt. For a summary of project related assets, visit our project website at https://self-rewarding-llm-training.github.io/.
@@ Appendix C Additional Experimental Results
C.1 Performance on Individual Test Datasets
In Figure 2, we showed the performance of SRT vs training with ground truth reward, averaged across 3 heldout test sets. Now we present their performance on each individual test set for the sake of completeness. Figures 10, 11, and 12 show the detailed results when we train on DAPO, MATH-12K, and AIME (1983-2023) respectively. Additionally, when training on MATH-12K, we also evaluate the intermediate checkpoints on the heldout set MATH-500, which is reported in Figure 11. Since MATH-500 contains 500 examples, calculating average@32 accuracy becomes expensive, and hence we could not use it as a test set for the other training setups.
@@ C.1 Performance on Individual Test Datasets
Figure 11: (Individual test set performance during training on MATH-12K) We record the average@32 accuracy during training Qwen2.5-Math-7B on MATH-12K, on three heldout test sets: AIME 2024, AIME 2025 and AMC. We also evaluate intermediate checkpoints on MATH-500 since we are training on MATH-12K (we could not do this for other training datasets due to a lack of computational resources). In all 4 heldout test sets, SRT results in similar performance gain as one would obtain from training with ground truth labels.
@@
C.2 Experimental Results Using a Different Base Model
In addition to Qwen2.5-Math-7B, we apply our algorithm with another LLM — namely Qwen3-14B-Base (Yang et al., 2025). We choose the base model since it has not gone through additional post-training on reasoning tasks, unlike the Qwen3-14B model. Additionally, this is a significantly larger model with a different pre-training, making it suitable for testing our algorithm’s effectiveness.
@@ C.2 Experimental Results Using a Different Base Model
Appendix D More on Enhancing SRT’s Performance
D.1 Training Dynamics of SRT on the Easy DAPO Subset
@@
Figure 16: (Training Dynamics of SRT on the Easy Subset of DAPO) We show the training dynamics of SRT on the easiest 1/3-rd of the DAPO dataset, chosen by ground truth pass rate of the base model. Compared to SRT on the entire DAPO dataset, SRT on the easier subset does not show any signs of reward hacking, even after taking 3 full passes over the training set.
@@
Figure 16 shows our results: SRT-training on the easiest subset does not show the same behavior as training on the full dataset: accuracy on the heldout set does not drop, and KL penalty, while being slightly higher than that of training with ground truth, is still significantly lower than SRT-training on the full dataset. We also see that model entropy does not explode, so the model keeps outputting reasonable responses instead of the degenerate ones resulting from full dataset training. The most intriguing observation is that regarding pseudo-reward (Figure 16, second from left): it very quickly gets very close to 1 and stabilizes around 0.9. This tends to suggest the model gets very little learning signal as the mean of the pseudo reward is already approximately 1, which is probably the reason it does not learn to reward hack within the same compute budget. We leave investigating this further for future work.
D.2 Training Qwen3-14B-Base on the Easy DAPO Subset
@@
Figure 17: (Qwen3-14B-Base Trained on the Easy DAPO Subset) We take the same easy subsets of DAPO from Section 5 and train a Qwen3-14B-Base model with SRT on it. We see the same behavior as Qwen2.5-Math-7B (Figure 8), that SRT on the easy subset does not exhibit performance collapse within the same compute budget.
@@
One of our most interesting observations is that simply using the easiest 1/3-rd of the DAPO dataset eliminates the performance collapse within our training budget (it can still happen if one trains more, though we do not observe it). We want to test whether this is still true for a different base model. To do so, we take the same easiest subset used in Section 5 (so the subset is determined using either the ground truth pass rate or the frequency of the majority answer of a Qwen2.5-Math-7B model) and train a Qwen3-14B-Base model with SRT on this subset. Figure 17 shows the result of our experiments: similar to Qwen2.5-Math-7B model, the Qwen3-14B-Base model also does not exhibit performance collapse within the same training budget.
Appendix E Behavior of a Model after Performance Collapse

### v2

@@ B.1 RL Algorithms
Sampling.
For all experiments, we kept the generation temperature to 1.0, top_k to -1, and top_p to 1 for rollouts generated during RL rollouts. Decoding temperature used for validation varies in different settings, see Appendix B.3 and D for more discussion. We cut off maximum prompt length at 1024 and maximum response length to 3072 (note: Qwen2.5-Math-7B models support a maximum context window of 4096).
@@ B.2 GPU Infrastructure
All experiments in this work were conducted using either a single node consisting of 8 NVIDIA H200 GPUs (141 GB of GPU memory per GPU) or a single node consisting of 4 NVIDIA GH200 GPUs (96 GB of GPU memory per GPU). All experiments can be replicated in single-node training, and we did not, in fact, utilize multinode training. In total, this work consumed $\sim$15000 GPU hours (including preliminary studies and failed runs). All the final results listed in this paper can be replicated within 2000 H200 GPU hours.
B.3 Details on Training Settings
We choose Family Relationships, Bitwise Arithmetic, and Knights & Knaves (Xie et al., 2024) tasks from Reasoning Gym (Stojanovski et al., 2025) for our experiments. Examples for each task is shown in Appendix K.
For the Bitwise Arithmetic task, Level refers to the difficulty parameter. The model was first trained with level 2 data for 950 steps, reaching 97% accuracy. We then used this initialized model to train with SRT on levels 3 and 4.
For the Family Relationships task, we trained a model to 99% accuracy on the level 4 dataset. Here, Level corresponds to the parameters min_family_size and max_family_size, both set to 4. We then applied SRT on level 5.
For the Knights & Knaves task, we varied only the n_people parameter as the difficulty control. We first trained a model with difficulty level 2 to 99% accuracy, and then used that checkpoint to further experiment with SRT, climbing from level 2 to 3, 3 to 5, 5 to 7 and 7 to 9. We only report levels 7 and 9 in the paper since they are the highest level difficulty among our experiments.
Across all multi-level experiments, we applied SRT progressively. For example, in Bitwise Arithmetic we trained on level 2 with ground truth supervision; then, starting from the level 2 checkpoint, we applied SRT on level 3; finally, we repeated SRT again on level 4 using the checkpoint from level 3. For comparing against SRT with fixed teacher, we use the same starting policy (Qwen3-4B-Base trained with ground-truth on the easiest difficulty level on each task) and generate the same number of rollouts per prompt using temperature 1.0 and perform majority voting among these rollouts to generate our pseudo-labels.
Default training hyperparameters for Reasoning Gym tasks.
For all Reasoning Gym experiments, we used the Qwen3-4B-Base model, with GRPO as the main algorithm. The learning rate was set to 1e-6 and the KL penalty to 0.0001. For all experiments, we used 32 rollouts per prompt for training and 16 rollouts for evaluation.
Default training hyperparameters for Math Datasets.
  • Qwen2.5-Math-7B: Learning rate $10^{-6}$, KL penalty coefficient $0.001$, decoding temperature for training and evaluation rollouts $1.0$, top-p $1.0$, and no top-k sampling.
  • Qwen3-14B-Base: Same default hyperparameter setting as Qwen2.5-Math-7B.
  • Llama-3.1-8B-Instruct: Learning rate $10^{-7}$, KL penalty coefficient $0.001$, decoding temperature for training is $1.0$ and evaluation is $0.0$, top-p $1.0$, and no top-k sampling. We subsample the Big-Math-RL-Verified dataset to only keep prompts that has average Llama-3.1-8B-Instruct pass rate between 0.3 and 0.7, since the model is unable to improve during training otherwise.
  • Deepseek-Math-7B-Instruct: Learning rate $10^{-7}$, KL penalty coefficient $0.001$, decoding temperature for training is $1.0$ and evaluation is $0.7$ (following official protocol), top-p $1.0$, and no top-k sampling.
B.4 Code and Dataset Release
@@
    The test datasets for math tasks compiled together: https://huggingface.co/datasets/ftajwar/srt_test_dataset
@@
    Individual training and evaluation datasets for various levels of the reasoning gym tasks (bitwise arithmetic, knights and knaves, family relationships) can be found here: https://huggingface.co/collections/ftajwar/self-rewarding-llm-training-6835218091832c3664176553.
  4. 4.
@@
Finally, our code is available at: https://github.com/tajwarfahim/srt. For a summary of project related assets, visit our project website at https://self-rewarding-llm-training.github.io/.
@@ Appendix C Additional Experimental Results
C.1 Qwen2.5-Math-7B
Here we compare the performance of training Qwen2.5-Math-7B with SRT and RL with ground truth on each individual test set for the sake of completeness. Figures 10, 11, and 12 show the detailed results when we train on DAPO, MATH-12K, and AIME (1983-2023) respectively. Additionally, when training on MATH-12K and DAPO, we also evaluate the intermediate checkpoints on the heldout set MATH-500, which is reported in Figure 11. Since MATH-500 contains 500 examples, calculating average@32 accuracy becomes expensive, and hence we could not use it as a test set for all our training setups.
@@
Figure 11: (Individual test set performance during training on MATH-12K) We record the average@32 accuracy during training Qwen2.5-Math-7B on MATH-12K, on three heldout test sets: AIME 2024, AIME 2025 and AMC. We also evaluate intermediate checkpoints on MATH-500 since we are training on MATH-12K (we could not do this for other training datasets due to a lack of computational resources). In all 4 heldout test sets, SRT results in similar performance gain as one would obtain from training with ground truth labels. However, performance collapses after 1200 RL steps, similar to our other observations.
@@
C.2 Qwen3-14B-Base
In addition to Qwen2.5-Math-7B, we apply our algorithm on another LLM — namely Qwen3-14B-Base (Yang et al., 2025). We choose the base model since it has not gone through additional post-training on reasoning tasks, unlike the Qwen3-14B model. Additionally, this is a significantly larger model with a different pre-training, making it suitable for testing our algorithm’s effectiveness.
@@ C.2 Qwen3-14B-Base
Appendix D Effect of Decoding Temperature (Qwen2.5-Math-7B)
(a) Math-500 evaluation accuracy during training, when decoded under various temperature (Qwen2.5-Math-7B).
(b) Absolute performance improvement after one epoch of training when decoded under various temperatures (Qwen2.5-Math-7B).
Figure 16: Our method (SRT) performs consistently regardless of decoding temperature for validation (Figure 16(a)). All experiments are run using Qwen2.5-Math-7B as the base model, trained on MATH-12K and tested on MATH-500. Notice that even though the performance is low initially at high temperature, at the later stages, they plateau around the same point. Figure 16(b) shows the absolute gain when decoded under different temperatures. Note that decoding with higher temperature might give the impression of a larger gain compared to low-temperature decoding. However, the evaluation curves during training resulting from SRT and RL with ground-truth look almost identical regardless of decoding temperature, which is one of our main observations in this work.
Appendix E Additional Self-Training Metrics
Figure 17: (Tracking format following success rate during SRT-training on Reasoning Gym) In order to track whether SRT is teaching reasoning strategies beyond just formatting the final answer correctly, we track two additional metrics throughout SRT-training: fraction of generations among all generations where the final answer is parseable and fraction of generations among those that are parseable where the final answer is correct. We see that due to training with RLVR on an easier level of difficulty, the starting policy can already format most generations correctly, and in the case of Knights & Knaves, fraction of correct responses keeps increasing even after fraction of properly formatted (and thus parseable) responses have saturated. This shows that the model learns reasoning strategies beyond formatting rules.
We are interested to know if SRT teaches actual reasoning strategies beyond just proper formatting rules necessary for extracting the final answer. To do so, we track two additional metrics throughout SRT-training: fraction of generations among all generations where the final answer is parseable and fraction of generations among those that are parseable where the final answer is correct. Figure 17 summarizes our findings on Reasoning Gym: We see that due to training with RLVR on an easier level of difficulty, the starting policy can already format most generations correctly, and in the case of Knights & Knaves, fraction of correct responses keeps increasing even after fraction of properly formatted (and thus parseable) responses have saturated.
Appendix F More Details on Baselines
Baseline Implementation.
For all three methods (SFT, DPO (Rafailov et al., 2024), ScPO (Prasad et al., 2024)), we sweep over three learning rates ($10^{-5}$, $10^{-6}$ and $10^{-7}$) and pick the checkpoint with the highest validation score. The best checkpoint with highest validation in the SFT stage has been used to initialize the DPO/ScPO training. We also train DPO with the above mentioned learning rates and picked the best score. For DPO and ScPO (Prasad et al., 2024), we used $\beta=0.1$, which we also found through sweep over (0.1, 0.3 and 0.5). Moreover, we add a negative log-likelihood loss with weight $1.0$ to the DPO and ScPO losses to stabilize them, similar to RPO (Pang et al., 2024b). We do not train for more than 1 epoch to prevent overfitting/unintentional unalignment (Tajwar et al., 2024; Razin et al., 2025) and fair comparison with SRT. ScPO was also performed for one round.
Train Dataset Method AMC/AIME MATH500
MATH SFT 0.18 0.75
ScPO 0.20 0.72
DPO 0.23 0.74
SRT (Ours) 0.32 0.80
DAPO SFT 0.18 0.75
ScPO 0.20 0.72
DPO 0.21 0.76
SRT (Ours) 0.31 0.75
Base Model Accuracy 0.15 0.42
Majority@32 Acc 0.20 0.79
Table 2: Comparison of different methods trained on either the MATH or DAPO dataset. Performance is evaluated on AMC/AIME24, 25 (average accuracy@32) and MATH500 (average accuracy@1). Notice that majority@32 accuracy scores are not directly comparable with the other accuracy metrics listed in the table.
Dataset Curation
For DPO and ScPO we labeled the most consistent response as the positive example and the least consistent response as the negative example for each question. Moreover, we only kept the instances where $w(x)$, the variance based weighing parameter (Prasad et al., 2024), was greater than 2.
Appendix G Detailed Experiment Results on Different Training Settings
G.1 GRPO vs RLOO
Figure 18: (GRPO vs RLOO comparison) We compare the behavior of SRT under two different RL optimization algorithm: GRPO vs RLOO. All experiments use a Qwen2.5-Math-7B model trained on DAPO, with the other hyperparameters being the default ones described in Appendix B.3. While SRT with GRPO seems to achieve higher performance than that of SRT employing RLOO, ultimately prolonged training using both algorithms lead to reward hacking and model collapse on all test datasets.
Figure 18 shows our experiment comparing how SRT behaves with different RL algorithms. In particular, we test two algorithms: GRPO and RLOO. While GRPO seems to achieve higher performance, both GRPO and RLOO training with SRT-reward leads to model collapse at similar number of steps — showing that the choice of the RL algo does not influence model collapse.
G.2 Different KL Penalty Coefficients
Figure 19: (SRT with different KL penalty coefficients) We compare the behavior of SRT with different KL penalty coefficients and report performance on all test datasets. All experiments here use a Qwen2.5-Math-7B model trained with RLOO on the DAPO dataset, with the hyperparameters other than KL penalty coefficient being the default ones described in Appendix B.3. Stronger KL penalty does not prevent or delay model collapse.
The most straightforward way of preventing reward hacking is to add a strong KL penalty to the training objective. In Figure 19, we explore this idea: to our surprise, we don’t find a higher KL penalty coefficient to delay or prevent model collapse. We attribute this to the reward hacking training signal being too strong to be overcome by the KL regularization.
G.3 Learning Rate
Figure 20: (SRT with different learning rate) We compare the behavior of SRT with different learning rate and report performance on all test datasets. All experiments here use a Qwen2.5-Math-7B model trained with RLOO on the DAPO dataset, with the hyperparameters other than learning rate being the default ones described in Appendix B.3. Lower learning rate tends to delay model collapse in our experiments — though we see performance decay on AIME 2024 and 2025 within our training budget, and hypothesize that training for longer with SRT, even with a smaller learning rate, will lead to complete model collapse as usual.
Another common hyperparameter to tune is the learning rate. To investigate the effect learning rate has on SRT, we finetune a Qwen2.5-Math-7B model with different learning rates using SRT on the DAPO dataset, with all other hyperparameters kept fixed at their default values described in Appendix B.3. Figure 20 shows our empirical findings: lowering learning rate seems to prevent model collapse within our training budget. However, we notice performance degradation on the harder AIME 2024 and 2025 datasets within our training budget, and hypothesize that prolonged training with SRT, even with a considerably lower learning rate, would still lead to model collapse. We could not study this in detail due to computational constraints, and leave this for future work.
G.4 Different Number of generations per prompt
Figure 21: (SRT with number of generations per prompt) We compare the behavior of SRT with different number of generations/rollouts per prompt and report performance on all test datasets. All experiments here use a Qwen2.5-Math-7B model trained with RLOO on the DAPO dataset, with the hyperparameters other than number of generations per prompt kept fixed at the default ones described in Appendix B.3. Surprisingly, lowering the number of generations per prompt seems to delay model collapse.
We observe the most surprising result among our hyperparameter tuning experiments when we vary the number of rollouts per prompt during training. Similarly as before, we finetune a Qwen2.5-Math-7B model with SRT on the DAPO dataset, with all hyperparameters (except number of rollouts per prompt) kept fixed at the their default values described in Appendix B.3. Figure 21 records our results: model collapse happens progressively later in the training run as we lower the number of rollouts per prompt. We hypothesize that generating fewer rollouts makes estimating the “true" majority voting label for a single prompt more noisy. This noisy estimation then makes hacking the reward signal harder as well, thereby delaying model collapse. We have not been able to study this phenomenon in more detail due to computational constraints, and leave studying this for future work.
G.5 Entropy Coefficient
Figure 22: (SRT with different entropy coefficient) We compare the behavior of SRT with different entropy coefficients and report performance on all test datasets. All experiments here use a Qwen2.5-Math-7B model trained with RLOO on the DAPO dataset, with the hyperparameters other than entropy coefficient kept fixed at the default ones described in Appendix B.3. Surprisingly, increasing the entropy coefficient hastens model collapse.
The final hyperparameter with which we experiment is adding an entropy loss to our regular RL objective. The modified RL objective is listed below:
$\mathcal{L}_{\text{entropy-augmented}}(\pi_{\theta})=\mathcal{L}_{\text{RL}}(\pi_{\theta})-\alpha\mathcal{H}(\pi_{\theta})$ (7)
where $\mathcal{L}_{\text{RL}}(\pi_{\theta})$ is the regular RL loss objective, $\alpha$ is the entropy coefficient, and $\mathcal{H}(\pi_{\theta})$ is the per-token entropy averaged across all tokens in all rollouts. A reasonable hypothesis is that adding entropy to the loss objective will prevent model collapse (Cheng et al., 2025) by discouraging the model to converge to one solution for every prompt. However, Figure 22 shows results in the contrary: adding an entropy term to the loss function and increasing the corresponding coefficient $\alpha$ accelerates model collapse. Upon inspection of the rollouts, we see that increasing $\alpha$ incentivizes the model to generate more random tokens in the rollouts, which maximizes entropy, followed by the same template final answer, which maximizes training (pseudo-)reward. This suggests that a better mechanism to prevent model diversity collapse is needed (Song et al., 2025a; Zhou et al., 2025), which we leave to future work to study.
Appendix H Test-Time Self-Improvement
H.1 SRT can be used for test-time training
Figure 23: (Test-Time Self-Training Performance) Given the test dataset $\mathcal{D}_{\text{test}}$, one can perform SRT on $\mathcal{D}_{\text{test}}$ before making predictions. Our results show that this improves the majority voting performance on $\mathcal{D}_{\text{test}}$ without access to ground truth labels. Notice that the y axis is showing majority@32 accuracy instead of average@32 accuracy, for a fairer comparison with the baseline.
We compare the test-time performance of majority voting after SRT test-time training as well as without any test-time training. Empirically, we observe (Fig 23) that test-time training via SRT provides relatively limited, yet noticeable, performance gains when measured under the maj@32 metric, compared to the popular majority voting baseline applied directly to outputs generated by the base model.
H.2 Why Doesn’t the Performance Collapse during Test-Time-Training?
[…]
Appendix I More on Model Collapse Resulting from Self-Training
As discussed before, the optimization objective in SRT can lead to significant initial improvements followed by eventual model collapse. Here, we explore complementary strategies to address model collapse and further enhance the performance achievable via self-training:
  1. (1)
  2. (2)
    An algorithmic strategy that delays model collapse by using pseudo-labels generated from a stable base model rather than from the continuously updated model.
  3. (3)
I.1 Early Stopping
In our algorithm, even a small labeled validation dataset can effectively identify the peak performance point during self-training, thereby mitigating model collapse. Figure 25 shows the progression of model performance, measured throughout training on the DAPO dataset and evaluated across several test sets. Crucially, we find that the peak performance consistently occurs around the same training step across different held-out datasets. The vertical line in Figure 25 marks early stopping using only 1% of DAPO as validation, with performance on other datasets remaining near-optimal.
I.2 Self-Training with Offline-Generated Labels
Figure 26: (Self-Training with Offline-Generated Labels) Performance comparison between SRT and its variant, where pseudo-labels are precomputed from a fixed base model checkpoint. The fixed teacher labels maintain training stability while achieving comparable performance to SRT, highlighting a limitation of the online labeling strategy. Note that training starts to collapse for MATH-12k dataset  1000 steps (which is roughtly  1.5 epoch).
The tendency toward model collapse arises because SRT prioritizes consistency over correctness, increasingly rewarding model agreement even if incorrect. A straightforward yet effective approach that delays (but not completely prevent) the model collapse is to get the pseudo-labels from a stable, previously fixed checkpoint, rather than leveraging labels from the evolving policy (but with the downside that the generated labels will not benefit from the improvement in majority voting as training progressed, observed in Figure 2). Here we generate pseudo-labels via majority voting rollouts from the Qwen2.5-Math-7B base model, store these offline-generated labels, and subsequently perform RL training against them. Figure 26 demonstrates that training with these offline-generated labels significantly stabilizes training while achieving performance comparable to SRT. Importantly, this indicates that the dynamic updating of pseudo-labels during training (online labeling) may not always confer substantial benefits, and instead can contribute to training instability. However, we noticed training collapse even with offline labeled datasets when trained for longer than one epoch for the MATH-12k dataset.
I.3 Self-Training with Curriculum Learning
Our third approach, curriculum learning, is motivated by the observation that the model experiences earlier collapse when training on the difficult DAPO dataset compared to the simpler MATH-12K dataset. The intuition is that, on a more challenging dataset, the model finds it easier to abandon its pretrained knowledge in favor of optimizing self-consistency rather than genuinely learning to solve the underlying task.
We leverage this hypothesis to implement a curriculum learning strategy (Bengio et al., 2009; Andrychowicz et al., 2017; Portelas et al., 2020; Florensa et al., 2017; Song et al., 2025b; Lee et al., 2025; Tajwar et al., 2025) by identifying the ‘easiest’ subset of the DAPO dataset. To be precise, we retain 1/3-rd of the easiest DAPO prompts selected according to two distinct metrics:
  1. (1)
    Pass rate of the base model, which utilizes ground-truth labels.
  2. (2)
Figure 27 shows that training on these easier subsets significantly delays the onset of reward hacking, allowing for continuous improvement even across multiple epochs. Remarkably, performance on these curriculum subsets reaches levels comparable to standard RL training with ground-truth labels on the entire DAPO dataset. More importantly, we did not observe training collapse even after  3 epochs of training on 1/3rd of "easy" DAPO dataset. These promising results suggest that curriculum strategies may further extend the benefits of SRT, which we leave as a future research direction.
I.4 Training Dynamics of SRT (Qwen2.5-Math-7B) on the Easy DAPO Subset
@@ I.3 Self-Training with Curriculum Learning
Figure 28: (Training Dynamics of SRT on the Easy Subset of DAPO, using Qwen2.5-Math-7B) We show the training dynamics of SRT on the easiest 1/3-rd of the DAPO dataset, chosen by ground truth pass rate of the base model. Compared to SRT on the entire DAPO dataset, SRT on the easier subset does not show any signs of reward hacking, even after taking 3 full passes over the training set.
@@
Figure 28 shows our results on Qwen2.5-Math-7B: SRT-training on the easiest subset does not show the same behavior as training on the full dataset: accuracy on the heldout set does not drop, and KL penalty, while being slightly higher than that of training with ground truth, is still significantly lower than SRT-training on the full dataset. We also see that model entropy does not explode, so the model keeps outputting reasonable responses instead of the degenerate ones resulting from full dataset training. The most intriguing observation is that regarding pseudo-reward (Figure 28, second from left): it very quickly gets very close to 1 and stabilizes around 0.9. This tends to suggest the model gets very little learning signal as the mean of the pseudo reward is already approximately 1, which is probably the reason it does not learn to reward hack within the same compute budget. We leave investigating this further for future work.
I.5 Training Qwen3-14B-Base on the Easy DAPO Subset
@@
Figure 29: (Qwen3-14B-Base Trained on the Easy DAPO Subset) We take the same easy subsets of DAPO used in Figure 28 and train a Qwen3-14B-Base model with SRT on it. We see the same behavior as Qwen2.5-Math-7B (Figure 27), that SRT on the easy subset does not exhibit performance collapse within the same compute budget.
@@
One of our most interesting observations is that simply using the easiest 1/3-rd of the DAPO dataset eliminates the performance collapse within our training budget (it can still happen if one trains more, though we do not observe it). We want to test whether this is still true for a different base model. To do so, we take the same easiest subset used in Figure 28 (so the subset is determined using either the ground truth pass rate or the frequency of the majority answer of a Qwen2.5-Math-7B model) and train a Qwen3-14B-Base model with SRT on this subset. Figure 29 shows the result of our experiments: similar to Qwen2.5-Math-7B model, the Qwen3-14B-Base model also does not exhibit performance collapse within the same training budget.
I.6 Generating Easy DAPO Subset using Qwen2.5-Math-1.5B
Figure 30: (Generating Easy DAPO Subset Using Qwen2.5-Math-1.5B) To see if the curriculum generation process is reproducible, we generate the easy 1/3-rd subset (by both majority voting frequency and pass rate) using a Qwen2.5-Math-1.5B model. This is in contrast with Figure 27 and Figure 29, where we used a Qwen2.5-Math-7B model for the easy subset generation. Furtheremore, we train the Qwen2.5-Math-1.5B on the easy subset as before, and make similar observations: training with SRT on a easy subset, even for 3 epochs, does not lead to performance collapse.
Next, we want to test if the easy subset generation process of our curriculum algorithm itself is reproducible using different base models. To do so, we generate the easy 1/3-rd subset using a Qwen2.5-Math-1.5B model by both majority voting frequency and pass rate. This is in contrast with our earlier sections, especially Figure 27 and Figure 29, where we used a Qwen2.5-Math-7B model for the subset generation. We also train the Qwen2.5-Math-1.5B model on the resulting easy subsets. Figure 30 shows our results — we see the same trend as our earlier result, i.e., training on the easy subsets, even for 3 epochs, does not lead to any performance collapse. Moreover, surprisingly, up to our training budget, SRT on the easy subset matches the performance of RL training with ground truth labels.
I.7 SRT training with online curriculum
Figure 31: (Dynamic oversampling then filtering delays model collapse within training budget) Here we experiment with the idea that dynamic oversampling then filtering prompts with low model self-consistency can prevent/delay model collapse by improving the training signal. All experiments are done using Qwen2.5-Math-7B trained on the DAPO dataset using hyperparameters described in Appendix B.3. For the dynamic sampling variant of SRT, we discard prompts where the majority voting answer does not occur at least 60% of all 32 generations for the prompt, and keep generating until we have at least 16 such prompts (the prompt batch size used for all baselines). We see that this dynamic oversampling then filtering variant of SRT can prevent model collapse within our training budget of 1000 RL steps, though prolonged training can still lead to model collapse in laters phases of the training.
Based on the insights we gained from the previous sections, the final strategy we experiments with to mitigate model collapse with SRT is dynamic oversampling then filtering the prompts, similar to the one introduced by DAPO (Yu et al., 2025). Concretely, let the prompt batch size be $B$. Then we employ the following routine: (1) Sample $B$ prompts from the training dataset, (2) Generate $K$ responses per prompt, calculate per prompt majority votes and the fraction of generations per prompt where the final answer is the same as the majority vote (i.e., self-consistency level), (3) Discard prompts where the majority voting fraction is smaller than a threshold, and repeat the prior steps until we have $B$ prompts and their corresponding generations, (4) Take one RL step on the generated dataset.
We investigate this idea with a Qwen2.5-Math-7B model trained on the DAPO dataset. We set the filtering threshold to be $0.6$ (so any prompt where the majority vote appears in less than 60% of all generations is discarded), and all other hyperparameters are kept fixed at the default ones from Appendix B.3. Figure 31 shows our empirical findings: we see that our intuition holds, and this dynamic oversampling and filtering procedure prevents model collapse within our training budget. We hypothesize that extended SRT-training would still lead to model collapse since the reward hacking solution is still an optimal solution under our training objective and it is still possible for the model to discover it during training. Despite this, our procedure of controlling the task distribution is effectively filtering away the harder prompts dynamically during training and employing an online curriculum, which results in delaying model collapse, just as we expect from our experiments using only offline curriculum for SRT-training in Appendix I.3.
Appendix J Detailed Experiment Results using non-Qwen models
To validate the efficacy of SRT on LLMs with different pre-training/post-training routine, we run additional experiments on two more models: Deepseek-Math-7B-Instruct (Shao et al., 2024) and Llama-3.1-8B-Instruct (Grattafiori et al., 2024). Our results are described below.
J.1 Deepseek-Math-7B-Instruct
We train Deepseek-Math-7B-Instruct on the MATH-12K dataset, and test on AIME 24, AIME 2025, AMC and MATH-500. For training, we use the same hyperparameters use used to train Qwen2.5-Math-7B-Instruct due to lack of compute for running a sweep over possible hyperparameters. We note that we did not find the recommended temperature or other sampling parameters for the Instruct model in (Shao et al., 2024), but their base models were evaluated with temperature 0.7, so we choose temperature 0.7, top-p 1.0 and no top-k sampling for our evaluations. Figure 32 shows our results: we see similar trends as our earlier experiments, where SRT initially matches performance gain from RL with ground truth, but then leads to performance collapse after prolonged training.
Figure 32: (Training Deepseek-Math-7B-Instruct on MATH-12K using SRT) We see similar trends for SRT-training on Deepseek-Math-7B-Instruct as we saw on our experiments with Qwen models: SRT initially matches performance gain obtained with RL training with ground truth, but faces performance collapse after prolonged training.
J.2 Llama-3.1-8B-Instruct
Figure 33: (Training Llama-3.1-8B-Instruct on Big-Math-RL-Verified with learning rate $10^{-7}$) Llama-3.1-8B-Instruct, when trained on a filter subset of the Big-Math dataset, shows significant gains on MATH-500 from both SRT and RL with ground truth. In fact, up to our training budget of 2K steps, both seem to improve performance at the same rate, from 52.6% to around 60%.
Figure 34: (Training Llama-3.1-8B-Instruct on Big-Math-RL-Verified with learning rate $3\times 10^{-7}$) Llama-3.1-8B-Instruct, when trained on a filter subset of the Big-Math dataset with a higher learning rate ($3\times 10^{-7}$ instead of $10^{-7}$ in Figure 33) demonstrates model collapse within the same training budget.
Llama-3.1-8B-Instruct showed no gains while being trained on DAPO or MATH-12K. This can be due to insufficient hyperparameter tuning or the model’s starting performance on these datasets not suitable for learning. So we chose the Big-Math dataset (Albalak et al., 2025), a dataset with over 250,000 math questions with verifiable answers. Moreover, this dataset has been constructed by filtering common evaluation datasets like MATH-500, making it suitable for our purposes. The primary benefit of using this dataset is that it comes with Llama-3.1-8B-Instruct pass rate, so we can easily ascertain the difficulty of each datapoint and aggregate a subset that can be suitable for training the Llama model. Specifically, we take the subset of Big-Math where Llama-3.1-8B-Instruct has pass rate between 0.3 and 0.7, to filter away too easy or too difficult questions. Next, we train the model on this subset using the same hyperparameters as we used for training the Qwen2.5-Math-7B-Instruct, except we lower the learning rate to $10^{-7}$ (from $10^{-6}$), as we see that leads to more stable learning curves. For evaluation, we use the same prompt template but temperature 0 (greedy decoding), to match the starting model’s performance reported on its model card, and report pass@1 accuracy.
Figure 33 shows our results: on MATH-500, Llama-3.1-8B-Instruct shows the same performance growth when trained via SRT or RL with ground truth, up to our training budget of 2000 steps. Performance growth is also significant, and training improves pass@1 accuracy from 52.6% to around 60%. Note that we do note report performance on the harder datasets like AIME, because the Llama model’s performance remain close to 0 throughtout training (with both objectives) on these datasets, signalling that they might be too hard for this model. We also ran an additional experiment using a higher learning rate of $3\times 10^{-7}$. Figure 34 shows our empirical findings: with the higher learning rate, Llama-3.1-8B-Instruct also start to show model collapse within our training budget.
Appendix K Example Tasks from Reasoning Gym
Example Tasks Task: Family Relationships (Level 4)
Question: John is married to Isabella. They have a child called Edward. Edward is married to Victoria. What is Isabella to Edward? Respond only with the word that describes their relationship. Answer: mother Task: Bitwise Arithmetic (Level 2)
Question: Please solve this problem. Assume there is arbitrary bit depth and that there are signed integers. If the answer is negative, reply as a negative value (e.g., 0x3-0x3), not the two’s-complement form. Reply only with the final hexadecimal value. $((0x3a24-0x24b8)+(0x1741\gg 0x3))$ Answer: 0x1854 Task: Knights and Knaves (Level 2)
Question: A very special island is inhabited only by sages and fools. Sages always tell the truth, and fools always lie. You meet 2 inhabitants: Zoey and Riley. Zoey commented, “Riley is a fool.” In Riley’s words: “Zoey is a sage or Riley is a sage.” So who is a sage and who is a fool? (Format your answer like: “Zoey is a sage/fool, and Riley is a sage/fool”) Answer: Zoey is a fool, and Riley is a sage.
Above of we see one example from each of the three Reasoning Gym tasks used in our work. The examples shown are of the lowest difficulties that we first train the model on using ground truth RL. We do self training on more difficult variants on the tasks. Difficulty can be changed by the modifying the either number of person or digits. In this work, we abstract it away by calling it "level".
Appendix L Behavior of a Model after Performance Collapse